# Best approximation of a three-variable function by sum of one-variable coordinate functions

### v1

@@ 3 Minimal Projection Cycles on a Finite Set
The relation (2.2) implies that, the error can be found by calculating a maximum over a finite set of the form $V=U_{x}\times U_{y}\times U_{z}$. In this section we will search for $X^{\prime}_{V,min}$.
@@
for a matrix $A$. Below, we will prove some necessary conditions for an ordered subset $p^{\prime}$ of $p$ to be an optimal projection cycle.
We start with a well-known algebraic definition and a lemma (See: Definition 1 and Proposition 7 in [22]).
@@
and call it the support of $v$. A solution $v_{0}$ of the homogeneous linear equation $Mv=0$ is called a minimal solution if the set $\operatorname{supp}(v_{0})$ is minimal, that is, there exist no other solution $v$ of the equation satisfying
$\operatorname{supp}(v)\varsubsetneqq\operatorname{supp}(v_{0})$
@@
Let $M$ be a matrix then a solution $v_{0}$ of the equation $Mv=0$ which $i_{1},i_{2},...,i_{t}$-th components are nonzero and all other components are zero, is minimal, then the column system consisting $i_{1},i_{2},...,i_{k}$-th columns of the matrix $M$ is a minimal linearly dependent column system.
@@
Let $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle and $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$. Then the vector, $\lambda=(\lambda_{1},...,\lambda_{k})$ is a minimal solution of the equation (3.1), where,
$\lambda_{i}=\begin{cases}\lambda^{\prime}_{i},\hskip 13.0pt\text{ if }i\in\{i_{1},i_{2},\dots,i_{t}\}\\ 0,\hskip 17.0pt\text{ otherwise }\end{cases}$ (3.2)
@@
Assume the contrary, suppose $\lambda$ is not a minimal solution of (3.1). Then, there exist a solution $\lambda^{0}$ of (3.1) such that,
@@
Thus, there exist a nonempty set $\{j_{1},j_{2},...,j_{s}\}\varsubsetneqq\{1,2,...,t\}$ such that the $i_{j_{1}},i_{j_{2}},...,i_{j_{s}}$-th components of $\lambda^{0}$ are nonzero and all other components are zero. And, this implies that
@@
The next theorem gives a necessary condition on minimality of a projection cycle.
@@
Let $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$ and the vector $\lambda$ be defined as in (3.2). Then, by the Lemma 3.2, $\lambda$ is a minimal solution of (3.1). Thus, by the Lemma 3.1 the proof follows. ∎
We can construct a system of equation with variables being the weights at points of $U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$ and equations being equations on the faces $x=0$, $x=1$, $y=0$, $y=1$, $z=0$ and $z=1$ correspondingly, as all other variables in (3.1) can be defined uniquely by the Remark 2.2. It is easy to verify that, this new system will be
@@
Reducing the matrix $M$ to row echelon form by using the fact that the rows $r_{1},\dots,r_{6}$ of the matrix satisfies the identity $r_{1}+r_{2}=r_{3}+r_{4}=r_{5}+r_{6},$ gives $M\sim M^{\prime}$, where $M^{\prime}$ is the matrix given below
$\begin{pmatrix}1&0&-1&-1&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1&0\\ 0&1&-1&-1&0&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1\\ 0&0&0&0&1&2&1&-1&-2&-1&-1&-1&0&1&1&0&-2&0&2&2&0&-2\\ 0&0&0&0&-1&1&2&1&-1&-2&0&-1&-1&0&1&1&-2&-2&0&2&2&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\end{pmatrix}$
By the Lemma 2.5 we know that there exist an optimal projection cycle on $U_{x}\times U_{y}\times U_{z}$ that has $\lambda(T_{1})>0$, $\lambda(T_{8})>0$, $\lambda(M_{i})\geq 0$, $1\leq i\leq 8$, $\lambda(F_{i})\geq 0$, $1\leq i\leq 6$, $\lambda(T_{i})\leq 0$, $2\leq i\leq 7$ and $\lambda(F_{i})\lambda(F_{j})=0$ if $i\neq j$ and $j-i$ is even. Thus, its enough to find all minimal solutions of $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0.$
Assume $v$ is a minimal solution of the equation $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0$. The Theorem 3.1 implies that the system of columns of $M^{\prime}$ corresponding to the nonzero components of $v$ is a minimal linear dependent system. It is not hard to verify that there are $123$ minimal column systems containing the first two columns and satisfying the above conditions. And these are the ones given below. Note that we give only the weights on the corner points and the points with positive weights. The remaining points and their weights can be reconstructed as explained above.
@@
1.1. $T_{1},T_{8},T_{2},T_{5}:\ \lambda(T_{1})=1,\ \lambda(T_{8})=1,\ \lambda(T_{2})=-1,\ \lambda(T_{5})=-1$
@@
[…]
@@
where $M_{5}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $M_{5,l_{i}}=\Pr_{l_{i}}M_{5}$, $i=3,4,5$, etc.

### v2

@@ 3 Minimal Projection Cycles on a Finite Set
for a matrix $A$. Below, we will prove some necessary conditions for an ordered subset $p^{\prime}$ of $p$ to be a minimal projection cycle.
We start with a well-known algebraic definition and a lemma (see: [37, Definition 1 and Proposition 7]).
@@
and call it the support of $v$. A solution $v_{0}$ of the homogeneous linear equation $Mv=0$ is called a minimal solution if the set $\operatorname{supp}(v_{0})$ is minimal, that is, there exists no other solution $v$ of the equation satisfying
$\operatorname{supp}(v)\varsubsetneqq\operatorname{supp}(v_{0}).$
@@
Let $M$ be a matrix then a solution $v_{0}$ of the equation $Mv=0,$ which $i_{1},i_{2},...,i_{t}$ -th components are nonzero and all other components are zero, is minimal, then the column system consisting of $i_{1},i_{2},...,i_{k}$ -th columns of the matrix $M$ is a minimal linearly dependent column system.
The next lemma connects minimal solutions and minimal projection cycles.
@@
Let $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle, and $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$. Then the vector, $\lambda=(\lambda_{1},...,\lambda_{k})$ is a minimal solution of the equation (3.1), where
$\lambda_{i}=\begin{cases}\lambda^{\prime}_{i},\hskip 13.0pt\text{ if }i\in\{i_{1},i_{2},\dots,i_{t}\}\\ 0,\hskip 17.0pt\text{ otherwise.}\end{cases}$ (3.2)
@@
Assume the contrary; suppose that $\lambda$ is not a minimal solution of (3.1). Then, there exists a solution $\lambda^{0}$ of (3.1) such that
@@
Thus, there exists a nonempty set $\{j_{1},j_{2},...,j_{s}\}\varsubsetneqq\{1,2,...,t\}$ such that the $i_{j_{1}},i_{j_{2}},...,i_{j_{s}}$-th components of $\lambda^{0}$ are nonzero and all other components are zero. And, this implies that
@@
The next theorem gives a necessary condition on the minimality of a projection cycle.
@@
Let $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$ and let the vector $\lambda$ be defined as in (3.2). Then, according to Lemma 3.2, $\lambda$ is a minimal solution of (3.1). Thus, by Lemma 3.1 the proof follows. ∎
Appendix: Finding all minimal projection cycles
The relation (2.14) implies that the error can be found by calculating a maximum over a finite set of the form $V=U_{x}\times U_{y}\times U_{z}$. Below, we will search for $V]^{(a)-(e)}_{min}-$the set of all minimal projection cycle-vectors on set $V$ satisfying all the propositions from the Theorem 2.1.
We can construct a system of equations with variables being the weights at the points of $U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$ and equations being equations on the faces $x=0$, $x=1$, $y=0$, $y=1$, $z=0$ and $z=1$ correspondingly, since the weights at all other points in $V$ can be defined uniquely as described in Remark 2.3. It is easy to verify that this system will be
@@
Reducing the matrix $M$ to the row echelon form by using the fact that the rows $r_{1},\dots,r_{6}$ of the matrix satisfy the identity $r_{1}+r_{2}=r_{3}+r_{4}=r_{5}+r_{6},$ gives $M\sim M^{\prime}$, where $M^{\prime}$ is the matrix given below
[…]
By the Lemma 2.5 we know that there exists an optimal projection cycle on $U_{x}\times U_{y}\times U_{z}$ that has $\lambda(T_{1})>0$, $\lambda(T_{8})>0$, $\lambda(M_{i})\geq 0$, $1\leq i\leq 8$, $\lambda(F_{i})\geq 0$, $1\leq i\leq 6$, $\lambda(T_{i})\leq 0$, $2\leq i\leq 7$ and $\lambda(F_{i})\lambda(F_{j})=0$ if $i\neq j$ and jij-i is even. Thus, it is sufficient to find the maximum over all minimal projection cycles on $V$ satisfying $\lambda(T_{1}),\lambda(T_{8})>0.$ Thus, by the Theorem 3.1, we can find all such projection cycles by finding all minimal solutions of the system $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0$ or in other words, finding all minimal linearly dependent column systems containing the first two columns.
It is not hard to verify with the help of a computer that there are $123$ minimal column systems containing the first two columns and corresponding to a minimal solution satisfying the above conditions. And corresponding projection cycle-vectors are the ones given below. Note that we give only the weights on the corner points and the points with positive weights. The remaining points and their weights can be reconstructed as explained above.
@@
$1.$ $T_{1},T_{8},T_{2},T_{5}:\ \lambda(T_{1})=1,\ \lambda(T_{8})=1,\ \lambda(T_{2})=-1,\ \lambda(T_{5})=-1$
@@
this is the function $f(x,y,z)=\begin{cases}xz,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ 0\leq x\leq\frac{1}{2},\ 0\leq z\leq\frac{1}{2},\\ \frac{z}{2}+\frac{(2x-1)y}{4},\ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ \frac{1}{2}\leq x\leq 1,\ 0\leq z\leq\frac{1}{2},\\ \frac{x}{2}+\frac{(2z-1)y}{4},\ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ 0\leq x\leq\frac{1}{2},\ \frac{1}{2}\leq z\leq 1,\\ \frac{1}{4}+\frac{(2x-1)y}{4}+\frac{(2z-1)y}{4},\ for\ \ \frac{1}{2}\leq x\leq 1,\ \frac{1}{2}\leq z\leq 1\end{cases}$:
@@
where $M_{5}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $M_{5,l_{i}}=\Pr_{l_{i}}M_{5}$, $i=3,4,5$.